The product of two consecutive even integers is 126 greater than three times the larger number. What are the two numbers?
Answer by CubeyThePenguin(3113)

consecutive even integers: x, (x+2)

x(x+2) = 126 + 3(x+2)
x^2 + 2x = 3x + 132
x^2 - x - 132 = 0
(x - 12)(x + 11) = 0

The integers are even, so the two numbers are 12 and 14.